tjmoney2424:

Simplify the following polynomial and write it in standard form: (5x2−3x+2)−(3x+5)+(4x2+10)\left(5x^2-3x+2\right)-\left(3x+5\right)+\left(4x^2+10\right)(5x 2 −3x+2)−(3x+5)+(4x 2 +10)

tjmoney2424:

(5x 2 −3x+2)−(3x+5)+(4x 2 +10)

xxDeppressionxx:

Simplify 2x2 - 3x - 5 25 - 4x2 Factoring 2x2 - 3x - 5 The first term is, 2x2 its coefficient is 2 . The middle term is, -3x its coefficient is -3 . The last term, "the constant", is -5 Step-1 : Multiply the coefficient of the first term by the constant 2 • -5 = -10 Step-2 : Find two factors of -10 whose sum equals the coefficient of the middle term, which is -3 . -10 + 1 = -9 -5 + 2 = -3 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -5 and 2 2x2 - 5x + 2x - 5 Step-4 : Add up the first 2 terms, pulling out like factors : x • (2x-5) Add up the last 2 terms, pulling out common factors : 1 • (2x-5) Step-5 : Add up the four terms of step 4 : (x+1) • (2x-5) Which is the desired factorization Factoring: 25-4x2 Theory : A difference of two perfect squares, A2 - B2 can be factored into (A+B) • (A-B) Proof : (A+B) • (A-B) = A2 - AB + BA - B2 = A2 - AB + AB - B2 = A2 - B2 Note : AB = BA is the commutative property of multiplication. Note : - AB + AB equals zero and is therefore eliminated from the expression. Check : 25 is the square of 5 Check : 4 is the square of 2 Check : x2 is the square of x1 Factorization is : (5 + 2x) • (5 - 2x) 3.3 Rewrite (5-2x) as (-1) • (2x-5) Cancel out (2x-5) which now appears on both sides of the fraction line. Pull out like factors : -x - 1 = -1 • (x + 1) -x - 1 ——— 2x + 5